OpenStudy (anonymous):

Can some one help me please ?? A soccer player throws the ball onto the field from the sideline. An equation modeling the balls flight is h=-16t^2+18t+2. In how many seconds will the ball reach the ground?? PLEASE HELP ME!

OpenStudy (anonymous):

hits the ground when h = 0 and t > 0 -16t^2 + 18t + 2 = 0 Divide through by -2 => 8t^2 - 9t - 1 = 0 Quadratic formula again... [ -(-9) +/- √(-9^2 - (4 * 8 * -1)) ] / (2 * 8) = [ 9 +/- √(81 + 32) ] / 16 = [ 9 +/- √(81 + 32) ] / 16 = [ 9 +/- √113 ] / 16 = -0.1 or 1.23 It can't be -0.1 so the answer is 1.23